OpenStudy (anonymous):

what will be the domain of (x+3)^1/3 plx tell me

OpenStudy (anonymous):

is 1

OpenStudy (anonymous):

tell me plx the answer of one question I have a quiz tomorrow

OpenStudy (anonymous):

how its domain is 1

OpenStudy (anonymous):

multiple 1/3 by x and 1/3 by 3 and x = 1

sam (.sam.):

Hold on, chrome crashed.

OpenStudy (anonymous):

ok

sam (.sam.):

First let \[y=(x+3)^{1/3}\] That's equivalent to \[\LARGE y=\sqrt[3]{x+3}\] To find the domain set y=0, \[0=x+3\] x=-3 Now we have "x greater than or equal to" (because it's positive values) -3 the values starts from -3 to infinity. These are real numbers also.

OpenStudy (anonymous):

so its domain would be all real number excepy -3?

sam (.sam.):

Nope, the domain will includes all the values greater than or equal to -3, which means \[\left\{ x \in \mathbb{R} : x \geq -3 \right\}\]

sam (.sam.):

include*

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

I have a one question plx solve it find the formulas for f+g,f-g,fg,f/g and domain of the function f(x)=2(x-1)^1/2 and g(x)=(x-1)^1/2

OpenStudy (anonymous):

^ 1/2 means under root

OpenStudy (anonymous):

just valus in bracket is in under root

sam (.sam.):

Its easy, for f+g, it means \[f(x)+g(x)\] \[\large 2(x-1)^{1/2} \color{red}{+} (x-1)^{1/2}\] -------------------------------------- For f-g, \[f(x)-g(x)\] \[\large 2(x-1)^{1/2} \color{red}{-} (x-1)^{1/2}\] -------------------------------------- For fg, \[f(x) \times g(x)\] \[\large 2(x-1)^{1/2} \color{red}{\times } (x-1)^{1/2}\]

OpenStudy (anonymous):

ok thank you and how can we find its domain I have problem in finding domain we will find separate of each?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have problem in finding domains

OpenStudy (anonymous):

what will be domain of 1-x

sam (.sam.):

That would be all real because there is no restrictions